### v2

Abstract
The stochastic Lanczos quadrature method has garnered significant attention recently. Upon examination of the error analyses given by Ubaru, Chen and Saad [SIAM J. Matrix Anal. Appl. (SIAMX), 38:1075–1099 (2017)] and Cortinovis and Kressner [Found. Comput. Math (FoCM), 22:875-903 (2021)], certain notable inconsistencies arise. It turns out that the former’s results are valid for cases with symmetric quadrature nodes and may not be adequate for many practical cases such as estimating log determinant of matrices. This paper analyzes probabilistic error bound of the stochastic Lanczos quadrature method for cases with asymmetric quadrature nodes. Besides, an optimized error allocation technique is employed to minimize the overall number of matrix vector multiplications required by the stochastic Lanczos quadrature method.
keywords:
log determinant approximation, trace estimation, stochastic Lanczos quadrature, asymmetric quadrature nodes
{AMS}
65C05, 65D32, 65F15, 65F60, 65G99, 65Y20, 68Q10, 68Q87
1 Background
The computation of log determinant arises in many proliferating applications, such as Gaussian process kernel learning [29], Bayesian interpolation [24], Kullback-Leibler divergence [20] and linear mixed models [5, 35, 36]. Since the data sets are continuously growing in the current era, a fast and scalable solver to calculate or approximate the log determinants of large matrices is in demand. The stochastic Lanczos quadrature (SLQ) method is a promising approach to approximate log determinants of large scale positive definite matrices. It might be first proposed by Bai, Fahey and Golub [2], while the first rigorous theoretical analysis was developed by Ubaru, Chen and Saad [33]. Recently, certain inconsistencies arise between their results and the work by Cortinovis and Kressner [6]. Since the SLQ method is receiving an increasing attention, a consistent and cogent result is necessary for further developing such a method [17]. This short note details these inconsistencies and develops a new error analysis for practical cases.
After a close examination of the SLQ method, it turns out that Ubaru, Chen and Saad’s result works for the cases when the quadrature nodes are symmetric [33]. Whereas Cortinovis and Kressner pointed out that quadrature nodes generated in the SLQ process may be asymmetric [6]. We propose Theorem 3.6 that guarantees the symmetric quadrature nodes under certain assumptions, but perhaps the asymmetric cases arise more often in practice. For instance, Fig. 1 illustrates that the quadrature nodes generated by SLQ are asymmetric based on the matrix nd3k from [7]. Another point worth noticing is that the SLQ method is based on a quadrature rule corresponding to a discontinuous measure defined by the spectrum of the underlying matrix (see (5) below). Unlike the continuous measure case, under such a discontinuous measure the quadrature error will remain unchanged when an affine transformation from the reference domain $\left[-1,1\right]$ to the physical domain $\left[\lambda_{\min},\lambda_{\max}\right]$ is applied (this will be detailed in Section 3.4). With consideration of these two points, this paper updates estimation on the query numbers and the Lanczos steps based on the asymmetric quadrature rule, and further develops a probabilistic relative error bound for log determinant estimation. The main results are presented in Theorem 3.16, Theorem 3.18 and Theorem 4.1.
(a) Asymmetric quadrature nodes on physical domain
(b) Asymmetric quadrature nodes on reference domain
(c) Asymmetric quadrature nodes on physical domain
(d) Asymmetric quadrature nodes on reference domain
Figure 1: Illustration of asymmetric quadrature nodes generated for the nd3k matrix from [7]. The initial normalized Rademacher vector $\boldsymbol{v}$ is chosen as $(1,\ldots,1,-1,\ldots,-1)^{T}/\sqrt{9000}\in\mathbb{R}^{9000}$, where there are equal numbers of 1s and -1s. Fig. 1a and Fig. 1c show quadrature nodes (eigenvalues) in the physical domain $\left[\theta_{1},\theta_{10}\right]$ and $\left[\theta_{1},\theta_{100}\right]$, while Fig. 1b and Fig. 1d depict the quadrature nodes that are transferred to the reference interval $[-1,1]$.
2 SLQ basics
For real symmetric positive definite matrix $A\in\mathbb{R}^{n\times n}$ with eigenvalues $\{\lambda_{j}\}_{j=1}^{n}$, the log determinant computation problem can be equivalently converted to the following terms [12, Lemma 11.6]
$\log\det A=\log\prod_{j=1}^{n}\lambda_{j}=\sum_{j=1}^{n}\log\lambda_{j}=\operatorname{tr}(\log(A)).$ (1)
There are various approaches [23] to estimate the trace term in equation (1) but this note would focus on the stochastic Lanczos quadrature (SLQ) method [2, 33]. For readers who are interested in other methods, please refer to papers regarding inducing points methods [27, 28], randomized low-rank approximation methods [32, 22], Chebyshev polynomial-based methods [15, 8], and methods that utilize variance reduction in trace estimation, such as Hutch++ and improved algorithms based on that [25, 26, 9].
The SLQ method utilizes the Girard-Hutchinson’s trace estimator [11, 19], where the trace of logarithm matrix in equation (1) can be approximated by an $N$-query Girard-Hutchinson’s trace estimator
$\operatorname{tr}(\log(A))\approx\operatorname{tr}_{N}(\log(A))=\frac{1}{N}\sum_{i=1}^{N}{\boldsymbol{z}^{(i)}}^{T}\log(A)\boldsymbol{z}^{(i)}=\frac{n}{N}\sum_{i=1}^{N}{\boldsymbol{v}^{(i)}}^{T}\log(A)\boldsymbol{v}^{(i)},$ (2)
where $\boldsymbol{z}^{(i)}$ is the $i^{th}$ query vector that follows the Rademacher distribution (i.e., every entry of the vector would either be +1 or -1 with probability 50%) and $\boldsymbol{v}^{(i)}=\boldsymbol{z}^{(i)}/\sqrt{n}$ is the corresponding unit vector. Based on the fact that $A$ is diagonalizable, its logarithm can be calculated by first taking the eigen-decomposition $A=Q\Lambda Q^{T}$ and then $\log(A)=Q\log(\Lambda)Q^{T}$. Note that $\log(\Lambda)$ is a diagonal matrix with entries $\log(\lambda_{j}),j=1,\ldots,n$. Let $\boldsymbol{\mu}^{(i)}=Q^{T}\boldsymbol{v}^{(i)}$, then equation (2) further reads
$\operatorname{tr}(\log(A))\approx\frac{n}{N}\sum_{i=1}^{N}{\boldsymbol{v}^{(i)}}^{T}\log(A)\boldsymbol{v}^{(i)}$ $=\frac{n}{N}\sum_{i=1}^{N}{\boldsymbol{v}^{(i)}}^{T}Q\log(\Lambda)Q^{T}\boldsymbol{v}^{(i)}$
$=\frac{n}{N}\sum_{i=1}^{N}\ {\boldsymbol{\mu}^{(i)}}^{T}\log(\Lambda)\boldsymbol{\mu}^{(i)}.$ (3)
[…]
$\frac{n}{N}\sum_{i=1}^{N}{\boldsymbol{\mu}^{(i)}}^{T}\log(\Lambda)\boldsymbol{\mu}^{(i)}$ $=\frac{n}{N}\sum_{i=1}^{N}\sum_{j=1}^{n}\log(\lambda_{j}^{(i)})\left[{\mu_{j}^{(i)}}\right]^{2}$ (4)
$=\frac{n}{N}\sum_{i=1}^{N}\int_{a}^{b}\log(t){\rm d}\mu^{(i)}(t)=\frac{n}{N}\sum_{i=1}^{N}I^{(i)},$
where $\mu_{j}^{(i)}$ is the $j^{th}$ element of $\boldsymbol{\mu}^{(i)}$ and $\mu^{(i)}(t)$ is the corresponding piecewise measure function of the $i^{th}$ integral
$\mu^{(i)}(t)=\left\{\begin{aligned} &\quad 0,&{\rm if}\ t<\lambda_{1}=a,\\ &\sum_{j=1}^{k-1}\left[{\mu_{j}^{(i)}}\right]^{2},&{\rm if}\ \lambda_{k-1}\leq t<\lambda_{k},k=2,...,n,\\ &\sum_{j=1}^{n}\left[{\mu_{j}^{(i)}}\right]^{2},&{\rm if}\ t\geq\lambda_{n}=b.\end{aligned}\right.$ (5)
According to the Gauss quadrature rule [12, Chapter 6.2], a Riemann Stieltjes integral $I^{(i)}$ can be approximated by an $(m+1)$-point Lanczos quadrature rule $I_{m}^{(i)}$ so the last term in equation (4) reads
$\frac{n}{N}\sum_{i=1}^{N}I^{(i)}\approx\frac{n}{N}\sum_{i=1}^{N}I_{m}^{(i)}=\frac{n}{N}\sum_{i=1}^{N}\sum_{k=1}^{m+1}\tau_{k}^{(i)}\log(\theta_{k}^{(i)}),$ (6)
where $m$ is the Lanczos step, and $\{\tau_{k}^{(i)}\}_{k=1}^{m+1}$, $\{\theta_{k}^{(i)}\}_{k=1}^{m+1}$ are the quadrature weights and nodes. According to [13], the weights $\{\tau_{k}^{(i)}\}_{k=1}^{m+1}$ are computed by the squares of the $1^{st}$ elements of the normalized eigenvectors $\{\boldsymbol{y}_{k}^{(i)}\}_{k=1}^{m+1}$ of $T_{m+1}^{(i)}$, and the nodes (also known as the Ritz values) $\{\theta_{k}^{(i)}\}_{k=1}^{m+1}$ are the corresponding eigenvalues $\{\lambda_{k}\}_{k=1}^{m+1}$ of $T_{m+1}^{(i)}$. Algorithm 1 [12, Section 4.1] outlines how to compute $m+1$ Lanczos quadrature nodes and weights given a real symmetric matrix and a randomly generated Rademacher vector.
Algorithm 1 Lanczos Algorithm for Lanczos Quadrature [12, Section 4.1]
Input: Real symmetric matrix $A\in\mathbb{R}^{n\times n}$, real vector $\boldsymbol{v}\in\mathbb{R}^{n\times 1}$, number of Lanczos iterations $m$.
Output: Quadrature nodes $\{\theta_{k}\}_{k=1}^{m+1}$ and quadrature weights $\{\tau_{k}\}_{k=1}^{m+1}$.
1:Obtain $\boldsymbol{v}_{1}=\boldsymbol{v}/\|\boldsymbol{v}\|_{2}$ and compute $\alpha_{1}={\boldsymbol{v}_{1}}^{T}A\boldsymbol{v}_{1}$.
2:Obtain $\tilde{\boldsymbol{v}}_{2}=A\boldsymbol{v}_{1}-\alpha_{1}\boldsymbol{v}_{1}$.
3:for $k=2$ to $m+1$ do
4:     Let $\beta_{k-1}=\|\tilde{\boldsymbol{v}}^{(k)}\|_{2}$.
5:     Obtain $\boldsymbol{v}_{k}=\tilde{\boldsymbol{v}}_{k}/\beta_{k-1}$.
6:     Compute $\alpha_{k}={\boldsymbol{v}_{k}}^{T}A\boldsymbol{v}_{k}$.
7:     Obtain $\tilde{\boldsymbol{v}}_{k+1}=A\boldsymbol{v}_{k}-\alpha_{k}\boldsymbol{v}_{k}-\beta_{k-1}\boldsymbol{v}_{k-1}$.
8:end for
9:Build a real symmetric tridiagonal matrix $T_{m+1}=\left[\begin{array}[]{ccccc}\alpha_{1}&\beta_{1}&0&\cdots&0\\ \beta_{1}&\alpha_{2}&\beta_{2}&\ddots&\vdots\\ 0&\beta_{2}&\ddots&\ddots&0\\ \vdots&\ddots&\ddots&\ddots&\beta_{m}\\ 0&\cdots&0&\beta_{m}&\alpha_{m+1}\end{array}\right].$
10:Calculate the eigenpairs of $T_{m+1}$: $[V,D]=\texttt{eig}(T_{m+1})$.
11:Obtain quadrature nodes $\theta_{k}$ as the diagonal entries of $D$, $k=1,\ldots,m+1$.
12:Compute quadrature weights $\tau_{k}=\left[\boldsymbol{e}_{1}^{T}\boldsymbol{y}_{k}\right]^{2}$, where $\boldsymbol{e}^{T}_{1}=(1,0,0,\ldots)$ and $\boldsymbol{y}_{k}$ is the $k^{th}$ column of $V$, $k=1,\ldots,m+1$.
13:return $\{\theta_{k}\}_{k=1}^{m+1}$ and $\{\tau_{k}\}_{k=1}^{m+1}$.
Algorithm 2 [33, Section 3] summarizes how the SLQ method approximates the log determinant for any SPD matrix.
Algorithm 2 Stochastic Lanczos Quadrature Method [2, 33]
Input: Real SPD matrix $A\in\mathbb{R}^{n\times n}$, number of Lanczos iterations $m$, and number of random vectors $N$.
Output: Approximation of log determinant of $A$: $\left[\log\det A\right]_{L}$.
1:for $i=1$ to $N$ do
2:     Randomly generate a Rademacher vector $\boldsymbol{z}^{(i)}\in\mathbb{R}^{n}$ and obtain $\boldsymbol{v}^{(i)}=\frac{\boldsymbol{z}^{(i)}}{\sqrt{n}}.$
3:     Apply $m$-step Algorithm 1 to $A$ with the initial vector $\boldsymbol{v}^{(i)}$ and obtain $\{\theta^{(i)}_{k}\}_{k=1}^{m+1}$ and $\{\tau_{k}^{(i)}\}_{k=1}^{m+1}$.
4:     $\Gamma\leftarrow\Gamma+\sum_{k=1}^{m+1}\tau_{k}^{(i)}\log(\theta_{k}^{(i)})$.
5:end for
6:return $\left[\log\det A\right]_{L}=\frac{n}{N}\Gamma$.
3 Error Analysis
According to [33, p. 9-10], the computational cost of stochastic Lanczos quadrature is $\mathcal{O}((\#nz(A)+nm^{2})N)$, where #$\#nz(A)$ denotes the number of non-zeros in symmetric matrix $A$ and $\mathcal{O}(nm^{2})$ represents the orthogonalization cost spent in the Lanczos iteration. Normally, the steps of Lanczos iteration $m$ and the number of Rademacher distributed vectors $N$ are expected to be much smaller than the matrix size $n$.
In the following of this paper, we denote the logarithm function by $f$ for simplicity.
3.1 Error for symmetric quadrature nodes
Ubaru, Chen and Saad [33, Theorem 4.2] study the absolute error between a Riemann Stieltjes integral and an $(m+1)$-point Lanczos quadrature.
Theorem 3.1.
[33, Theorem 4.2] Let $f$ be analytic in $\left[-1,1\right]$ and analytically continuable in the open Bernstein ellipse $E_{\rho}$ with foci ±$\pm 1$ and elliptical radius $\rho>1$. Let $M_{\rho}$ be the maximum of $|g(t)|$ on $E_{\rho}$. Then the $(m+1)$-point Lanczos quadrature approximation satisfies
$|I-I_{m}|\leq\frac{4M_{\rho}}{1-\rho^{-2}}\rho^{-2m-2}.$ (7)
Proof 3.2.
See [33, Section 4.1]. It should be pointed out that the result of Theorem 3.1 assumes that the quadrature nodes are symmetric.
As a direct consequence, they give a probabilistic absolute error bound for SLQ’s log determinant approximation based on symmetric quadrature rules.
Corollary 3.3.
[33, Corollary 4.5] Let $A\in\mathbb{R}^{n\times n}$ be a real symmetric positive definite matrix with eigenvalues in $[\lambda_{\min},\lambda_{\max}]$, and $\kappa=\lambda_{\max}/\lambda_{\min}$ be the condition number. Given $\epsilon,\eta\in(0,1)$, if the Lanczos iteration step $m$ and the query number $N$ satisfy
$m\geq(\sqrt{3\kappa}/4)\cdot\log(K/\epsilon),$
and
$N\geq(24/\epsilon^{2})\cdot\left(\log(1+\kappa)\right)^{2}\cdot\log(2/\eta),$
where
$K=\frac{5\kappa\log(2(\kappa+1))}{\sqrt{2\kappa+1}},$
then
$\mathbb{P}\Big{\{}\left|\log\det A-\left[\log\det A\right]_{L}\right|\leq\epsilon n\Big{\}}\geq 1-\eta.$
Proof 3.4.
See [33, Section 4.3].
3.2 Existence of symmetric quadrature nodes
As previously mentioned, Theorem 3.1 and Corollary 3.3 are valid when the Gauss quadrature nodes are symmetric, but when do these nodes preserve this property? In order to answer this question, we propose the following Theorem 3.6 with the aid of Definition 3.5.
Definition 3.5.
A vector $\boldsymbol{v}$ is said to have symmetric absolute values if
$|v_{i}|=|v_{n+1-i}|,i=1,\cdots,\lfloor\frac{n}{2}\rfloor.$
Theorem 3.6.
Let $A=Q\Lambda Q^{T}\in\mathbb{R}^{n\times n}$ be a symmetric positive definite matrix with eigenvalues $\lambda_{1},\cdots,\lambda_{n}$, which are symmetric about $\bar{\lambda}$, and $\boldsymbol{v}\in\mathbb{R}^{n}$ be an initial vector for Lanczos iteration. If vector $\boldsymbol{\mu}_{1}=Q^{T}\boldsymbol{v}$ has symmetric absolute values, then the distribution of $(m+1)$ Ritz values generated by the $m$-step Lanczos iteration is symmetric about $\bar{\lambda}$, where $m<n$.
Note that it is not strict to assume $A$ having symmetric eigenvalue distribution. For instance, certain matrices with property $\mathscr{A}$ [14, Definition 9-2.1] meet the condition.
Definition 3.7 (Definition 9-2.1 in [14]).
Let $A\in\mathbb{C}^{n\times n}$ be partitioned into the form
$A=\begin{bmatrix}A_{1,1}&\cdots&A_{1,q}\\ \vdots&&\vdots\\ A_{q,1}&\cdots&A_{q,q}\end{bmatrix},$ (8)
[…]
In other words, the vertices of matrices that share property $\mathscr{A}$ can be partitioned in two sets so that no edges exist in each set. Such matrices can be permuted into
$\begin{bmatrix}D_{1}&F\\ E&D_{2}\end{bmatrix},$ (9)
where $D_{1}$ and $D_{2}$ are diagonal matrices [31, Section 4.2.5]. Jacobi iteration matrix $C\in\mathbb{C}^{(m+n)\times(m+n)}$ is a special kind of (9),
$C=\begin{bmatrix}O&B\\ B^{H}&O\end{bmatrix},$ (10)
where $O$ denotes matrix of zeros and $B\in\mathbb{C}^{m\times n}$. Such matrices have symmetric eigenvalue distribution about 00, proved by Theorem A.1 in Appendix A [21, 3]. Here we give a more concise theorem that proves the symmetry of eigenvalues when $m=n$.
Theorem 3.8.
Let $B\in\mathbb{C}^{n\times n}$ and $B=U\Sigma V^{H}$ be its singular value decomposition, where $\Sigma=\mathrm{diag}(\sigma_{1},\cdots,\sigma_{n})$, $U=\begin{bmatrix}\boldsymbol{u}_{1},\cdots,\boldsymbol{u}_{n}\end{bmatrix}$ and $V=\begin{bmatrix}\boldsymbol{v}_{1},\cdots,\boldsymbol{v}_{n}\end{bmatrix}$. Then the matrix
$C=\begin{bmatrix}O&B\\ B^{H}&O\end{bmatrix}\in\mathbb{C}^{2n\times 2n}$
has a symmetric eigenvalue distribution. Precisely, $\left(\sigma_{i},\begin{bmatrix}\boldsymbol{u}_{i}\\ \boldsymbol{v}_{i}\end{bmatrix}\right)$ and $\left(-\sigma_{i},\begin{bmatrix}\boldsymbol{u}_{i}\\ -\boldsymbol{v}_{i}\end{bmatrix}\right),\\ i=1,\cdots,n$ are the eigenpairs of $C$.
Proof 3.9.
This result is more detailed than Proposition 4.12 in [31], which exactly descripts all the eigenpairs of $C$ in terms of the singular value decomposition. Based on $B=U\Sigma V^{H}$, it is trivial to obtain $BV=U\Sigma$ and $B^{H}U=V\Sigma$, i.e., $B\boldsymbol{v}_{i}=\sigma_{i}\boldsymbol{u}_{i}$ and $B^{H}\boldsymbol{u}_{i}=\sigma_{i}\boldsymbol{v}_{i}$, $i=1,\cdots,n$. Then
$C\begin{bmatrix}\boldsymbol{u}_{i}\\ \boldsymbol{v}_{i}\end{bmatrix}=\begin{bmatrix}O&B\\ B^{H}&O\end{bmatrix}\begin{bmatrix}\boldsymbol{u}_{i}\\ \boldsymbol{v}_{i}\end{bmatrix}=\begin{bmatrix}B\boldsymbol{v}_{i}\\ B^{H}\boldsymbol{u}_{i}\end{bmatrix}=\begin{bmatrix}\sigma_{i}\boldsymbol{u}_{i}\\ \sigma_{i}\boldsymbol{v}_{i}\end{bmatrix}=\sigma_{i}\begin{bmatrix}\boldsymbol{u}_{i}\\ \boldsymbol{v}_{i}\end{bmatrix},$
$C\begin{bmatrix}\boldsymbol{u}_{i}\\ -\boldsymbol{v}_{i}\end{bmatrix}=\begin{bmatrix}O&B\\ B^{H}&O\end{bmatrix}\begin{bmatrix}\boldsymbol{u}_{i}\\ -\boldsymbol{v}_{i}\end{bmatrix}=\begin{bmatrix}-B\boldsymbol{v}_{i}\\ B^{H}\boldsymbol{u}_{i}\end{bmatrix}=\begin{bmatrix}-\sigma_{i}\boldsymbol{u}_{i}\\ \sigma_{i}\boldsymbol{v}_{i}\end{bmatrix}=-\sigma_{i}\begin{bmatrix}\boldsymbol{u}_{i}\\ -\boldsymbol{v}_{i}\end{bmatrix},$
which proves the result.
Now we have checked the existence of matrices with symmetric eigenvalue distribution. Furthermore, when the conditions in Theorem 3.6 hold, the Jacobi matrix $T_{m+1}$ generated by $m$-step Lanczos iteration has constant diagonal entries.
Lemma 3.10.
Let $A=Q\Lambda Q^{T}\in\mathbb{R}^{n\times n}$ be a symmetric positive definite matrix with eigenvalues $\lambda_{1},\cdots,\lambda_{n}$, which are symmetric about $\bar{\lambda}$, and $\boldsymbol{v}\in\mathbb{R}^{n}$ be an initial vector for Lanczos iteration. If vector $\boldsymbol{\mu}_{1}=Q^{T}\boldsymbol{v}$ has symmetric absolute values, then the Jacobi matrix $T_{m+1}$ generated by $m$-step Lanczos iteration has constant diagonal entries $\bar{\lambda}$, where $m<n$.
Proof 3.11.
The proof is divided into two cases: even $n$ and odd $n$. We first study the even case and then extend the conclusion to the odd one. For $m$-step Lanczos iteration with initial vector $\boldsymbol{v}_{1}=\boldsymbol{v}$, the values $\alpha_{k}$ are computed as
$\alpha_{k}=\boldsymbol{v}_{k}^{T}A\boldsymbol{v}_{k}=\boldsymbol{v}_{k}^{T}Q\Lambda Q^{T}\boldsymbol{v}_{k}=\boldsymbol{\mu}_{k}^{T}\Lambda\boldsymbol{\mu}_{k},k=1,\cdots,m+1,$
where normalized vectors $\boldsymbol{v}_{k}=\boldsymbol{\tilde{v}}_{k}/\|\boldsymbol{\tilde{v}}_{k}\|$ and $\boldsymbol{\mu}_{k}=Q^{T}\boldsymbol{v}_{k},k=2,\cdots,m+1$. There exists a three-term recurrence between vectors $\boldsymbol{\tilde{v}}_{k}$ [12],
$\boldsymbol{\tilde{v}}_{k}=A\boldsymbol{v}_{k-1}-\alpha_{k-1}\boldsymbol{v}_{k-1}-\beta_{k-2}\boldsymbol{v}_{k-2},k=3,\cdots,m+1,$
$\boldsymbol{\tilde{v}}_{2}=A\boldsymbol{v}_{1}-\alpha_{1}\boldsymbol{v}_{1}$. For $k=1$, the computation of $\alpha_{1}$ reads
$\alpha_{1}$ $=\boldsymbol{v}_{1}^{T}A\boldsymbol{v}_{1}=\boldsymbol{v}_{1}^{T}Q\Lambda Q^{T}\boldsymbol{v}_{1}=\boldsymbol{\mu}_{1}^{T}\Lambda\boldsymbol{\mu}_{1}=\sum_{i=1}^{n}\mu_{1,i}^{2}\lambda_{i}$
$=(\lambda_{1}+\lambda_{n})\mu_{1,1}^{2}+(\lambda_{2}+\lambda_{n-1})\mu_{1,2}^{2}+\cdots+(\lambda_{\frac{n}{2}}+\lambda_{\frac{n}{2}+1})\mu_{1,\frac{n}{2}}^{2}$
$=2\bar{\lambda}\sum_{i=1}^{n/2}\mu_{1,i}^{2}=\bar{\lambda}.$
The last equal sign results from the assumption that $\boldsymbol{\mu}_{1}$ has symmetric absolute values and $\|\boldsymbol{\mu}_{1}\|=1$, indicating $\sum_{i=1}^{n/2}\mu_{1,i}^{2}=1/2$. For $k=2$, $\tilde{\boldsymbol{\mu}}_{2}$ reads
$\tilde{\boldsymbol{\mu}}_{2}=Q^{T}\boldsymbol{\tilde{v}}_{2}=Q^{T}Q\Lambda Q^{T}\boldsymbol{v}_{1}-\alpha_{1}Q^{T}\boldsymbol{v}_{1}=(\Lambda-\bar{\lambda}I)\boldsymbol{\mu}_{1}.$
Let $\tilde{\lambda}_{i}=\lambda_{i}-\bar{\lambda},i=1,\cdots,n,$ and
$\boldsymbol{\mu}_{1}=\begin{bmatrix}\mathrm{sgn}(\mu_{1,1})|\mu_{1,1}|\\ \vdots\\ \mathrm{sgn}(\mu_{1,n})|\mu_{1,n}|\end{bmatrix},$
then
$\tilde{\boldsymbol{\mu}}_{2}=\begin{bmatrix}\mathrm{sgn}(\mu_{1,1})\tilde{\lambda_{1}}|\mu_{1,1}|\\ \mathrm{sgn}(\mu_{1,2})\tilde{\lambda_{2}}|\mu_{1,2}|\\ \vdots\\ -\mathrm{sgn}(\mu_{1,n-1})\tilde{\lambda_{2}}|\mu_{1,n-1}|\\ -\mathrm{sgn}(\mu_{1,n})\tilde{\lambda_{1}}|\mu_{1,n}|\end{bmatrix},$
since $\tilde{\lambda}_{i}=-\tilde{\lambda}_{n+1-i},i=1,\cdots,n/2$. It is obvious that $\tilde{\boldsymbol{\mu}}_{2}$ have symmetric absolute values. For its normalized vector $\boldsymbol{\mu}_{2}=\tilde{\boldsymbol{\mu}}_{2}/\|\tilde{\boldsymbol{\mu}}_{2}\|$, one can deduce $\sum_{i=1}^{n/2}\mu_{2,i}^{2}=1/2$ and $\alpha_{2}=\bar{\lambda}$.
Let us further check $\tilde{\boldsymbol{\mu}}_{3}$.
$\tilde{\boldsymbol{\mu}}_{3}=Q^{T}\boldsymbol{\tilde{v}}_{3}$ $=Q^{T}Q\Lambda Q^{T}\boldsymbol{v}_{2}-\alpha_{2}Q^{T}\boldsymbol{v}_{2}-\beta_{1}Q^{T}\boldsymbol{v}_{1}=(\Lambda-\bar{\lambda}I)\boldsymbol{\mu}_{2}-\beta_{1}\boldsymbol{\mu}_{1}$
$=\begin{bmatrix}\mathrm{sgn}(\mu_{1,1})(\tilde{\lambda}_{1}^{2}/\|\tilde{\boldsymbol{\mu}}_{2}\|-\beta_{1})|\mu_{1,1}|\\ \mathrm{sgn}(\mu_{1,2})(\tilde{\lambda}_{2}^{2}/\|\tilde{\boldsymbol{\mu}}_{2}\|-\beta_{1})|\mu_{1,2}|\\ \vdots\\ \mathrm{sgn}(\mu_{1,n-1})((-(-\tilde{\lambda}_{2}^{2})/\|\tilde{\boldsymbol{\mu}}_{2}\|-\beta_{1})|\mu_{1,n-1}|\\ \mathrm{sgn}(\mu_{1,n})((-(-\tilde{\lambda}_{1}^{2})/\|\tilde{\boldsymbol{\mu}}_{2}\|-\beta_{1})|\mu_{1,n}|\end{bmatrix},$
which proves $\tilde{\boldsymbol{\mu}}_{3}$ have symmetric absolute values and leads to $\alpha_{3}=\bar{\lambda}$. Further steps of $k=4,\cdots,m+1$ will not be shown here since the recurrence relationship is exactly the same as how $\tilde{\boldsymbol{\mu}}_{3}$ is computed. The proof of even case ends.
When $n$ is odd, for any $k=1,\cdots,m+1$, $\alpha_{k}$ reads
$\alpha_{k}=\boldsymbol{\mu}_{k}^{T}\Lambda\boldsymbol{\mu}_{k}=\sum_{i=1}^{\lfloor\frac{n}{2}\rfloor}\mu_{k,i}^{2}\lambda_{i}+\mu_{k,\lfloor\frac{n}{2}\rfloor+1}^{2}\bar{\lambda}+\sum_{i=\lfloor\frac{n}{2}\rfloor+2}^{n}\mu_{k,i}^{2}\lambda_{i}.$
Here the two summations are still equal and $\|\boldsymbol{\mu}_{k}\|=1$, deducing
$\alpha_{k}$ $=(\lambda_{1}+\lambda_{n})\mu_{k,1}^{2}+(\lambda_{2}+\lambda_{n-1})\mu_{k,2}^{2}+\cdots+(\lambda_{\lfloor\frac{n}{2}\rfloor}+\lambda_{\lfloor\frac{n}{2}\rfloor+2})\mu_{k,\lfloor\frac{n}{2}\rfloor}^{2}$
$\quad+\bar{\lambda}\mu_{k,\lfloor\frac{n}{2}\rfloor+1}^{2}$
$=2\bar{\lambda}(1-\mu_{k,\lfloor\frac{n}{2}\rfloor+1}^{2})/2+\bar{\lambda}\mu_{k,\lfloor\frac{n}{2}\rfloor+1}^{2}$
$=\bar{\lambda},$
so $T_{m+1}$ has constant diagonal entries $\bar{\lambda}$.
It is trivial that every symmetric tridiagonal matrix with constant diagonal elements $\bar{\lambda}$ can be permuted into
$\begin{bmatrix}\bar{\lambda}I_{m}&B\\ B^{H}&\bar{\lambda}I_{n}\end{bmatrix}.$ (11)
Thus, based on Corollary A.3, the Jacobi matrix $T_{m+1}$ generated by $m$-step Lanczos iteration have symmetric eigenvalue distribution, indicating that the $(m+1)$ quadrature nodes are symmetric about $\bar{\lambda}$. In addition, one may find several tests on Theorem 3.6 in Appendix B.
3.3 Error for asymmetric quadrature nodes
Though Theorem 3.6 gives sufficient conditions for symmetric quadrature nodes, the conditions are not always satisfied in practice. As shown in Section 1, Fig. 1 demonstrates that Lanczos quadrature nodes might be asymmetric. Therefore, Theorem 3.12 derives an error estimation for the Lanczos quadrature with asymmetric quadrature nodes.
Theorem 3.12.
[6, Corollary 3] Let $g$ be analytic in $\left[-1,1\right]$ and analytically continuable in the open Bernstein ellipse $E_{\rho}$ with foci ±$\pm 1$ and elliptical radius $\rho>1$. Let $M_{\rho}$ be the maximum of $|g(t)|$ on $E_{\rho}$. Then the $(m+1)$-point Lanczos quadrature approximation with asymmetric quadrature nodes satisfies
$|I-I_{m}|\leq\frac{4M_{\rho}}{1-\rho^{-1}}\rho^{-2m-2}.$ (12)
Proof 3.13.
The proofs of Theorem 3.12 and Theorem 3.1 differ by a single step, namely the calculation of an infinite geometric progression
$|I-I_{m}|\leq\sum_{j=2m+2}^{\infty}\frac{4M_{\rho}}{\rho^{j}}.$ (13)
As stated that the Lanczos quadrature rule is not always symmetric for the measure (5) [6, Theorem 3], the progression reads
$\sum_{j=2m+2}^{\infty}\frac{4M_{\rho}}{\rho^{j}}=\frac{4M_{\rho}}{\rho^{2m+2}}\cdot\frac{1}{1-\rho^{-1}}=\frac{4M_{\rho}}{1-\rho^{-1}}\rho^{-2m-2}.$ (14)
Remark 3.14.
Theorem 3.12 is in fact [6, Corollary 3] but with a redefinition of the parameter $m$. In [6, Algorithm 1], $m$ corresponds to $m$ quadrature nodes. Whereas, in Theorem 3.12, $m$ corresponds to $(m+1)$ quadrature nodes, so does [33].
3.4 Error under affine transform
Theorem 3.12 assumes that function $g$ is analytic in the reference domain $\left[-1,1\right]$ and then gives an error estimation in the physical domain $\left[\lambda_{\min},\lambda_{\max}\right]$. In order to conduct analysis back on the physical interval, a proper affine transform should be included. The SLQ method is based on a quadrature rule corresponding to a discontinuous measure (5). Unlike the continuous measure case, under such a discontinuous measure the quadrature error will remain unchanged when using an affine transformation from the reference domain $\left[-1,1\right]$ to the physical domain $\left[\lambda_{\min},\lambda_{\max}\right]$ [16]. Precisely, consider the integrals
[…]
and
[…]
where $\mu^{(i)}$ are defined in (5), $h$ is an affine transform that shifts the interval $[-1,1]$ to $[\lambda_{\min},\lambda_{\max}]$, $g=f\circ h$, and $g$ satisfies the assumptions in Theorem 3.12. These integrals are in fact the finite sums of the form
[…]
$=\sum_{j=1}^{n}f(\lambda_{j})\left[\mu^{(i)}(\lambda_{j})-\mu^{(i)}(\lambda_{j}-)\right]$
[…]
where $\mu^{(i)}(t-)$ represents the left limit of the function $\mu^{(i)}$ on $t$. One may proceed another computation with the delta function
[…]
$=\int_{-1}^{1}f(h(t))\sum_{j=1}^{n}\left[\mu_{j}^{(i)}\right]^{2}\delta(h(t)-\lambda_{j})\left(\frac{\lambda_{\max}-\lambda_{\min}}{2}\right)dt$
$=\left(\frac{\lambda_{\max}-\lambda_{\min}}{2}\right)\sum_{t_{j}:h(t_{j})=\lambda_{j}}\left[\mu_{j}^{(i)}\right]^{2}f(h(t_{j}))$
$=\left(\frac{\lambda_{\max}-\lambda_{\min}}{2}\right)\sum_{j=1}^{n}\left[\mu_{j}^{(i)}\right]^{2}f(\lambda_{j})$
$=\left(\frac{\lambda_{\max}-\lambda_{\min}}{2}\right)I^{(i)},$
which gives rise to an extra scalar $(\lambda_{\max}-\lambda_{\min})/2$. Such a derivation assumes the validity of the chain rule of measures
$d\mu^{(i)}(h(t))$ $=\frac{d\mu^{(i)}}{dt}(h(t))\frac{dh}{dt}(t)dt$ (15)
$=\left(\frac{\lambda_{\max}-\lambda_{\min}}{2}\right)\sum_{j=1}^{n}\left[\mu_{j}^{(i)}\right]^{2}\delta(t-h^{-1}(\lambda_{j}))dt$
$=\left(\frac{\lambda_{\max}-\lambda_{\min}}{2}\right)d\mu^{(i)}(h(t)),$
where $d\mu^{(i)}(h(t))$ is the measure with distribution function $\mu^{(i)}(h(t))$ and $dt$ is the Lebesgue measure. However, this will lead to contradiction if $\lambda_{\max}-\lambda_{\min}\neq 2$. In fact, the chain rule of measures
$\frac{d\nu}{d\lambda}=\frac{d\nu}{d\mu}\frac{d\mu}{d\lambda}$
is guaranteed by some assumptions on $\nu,\mu$ and $\lambda$. One such assumption is given in [10].
Proposition 3.15.
[10, Proposition 3.9] Suppose $\nu,\mu,\lambda$ are $\sigma$-finite measures on $(X,\mathcal{M})$ such that
$\nu(E)=0\mbox{ for all }\mu(E)=0,\mbox{ and }\mu(E)=0\mbox{ for all }\lambda(E)=0,E\in\mathcal{M},$ (16)
then the Radon-Nikodym derivatives satisfy the chain rule
$\frac{d\nu}{d\lambda}=\frac{d\nu}{d\mu}\frac{d\mu}{d\lambda},\lambda-a.e.$
The measure in (15) does not satisfy assumption (16) since $d\mu^{(i)}(h(t))$ assigns positive measure $\left[\mu_{j}^{(i)}\right]^{2}$ on single point $h^{-1}(\lambda_{j})$, while the Lebesgue measure $dt$ assigns zero on any single point. This may be the reason that (15) fails.
3.5 Analyses for SLQ with asymmetric quadrature nodes
In light of the framework of Corollary 3.3, we derive an analysis for the SLQ method in Theorem 3.16.
Theorem 3.16.
Let $A\in\mathbb{R}^{n\times n}$ be a real symmetric positive definite matrix with eigenvalues in $[\lambda_{\min},\lambda_{\max}]$, and $\kappa=\lambda_{\max}/\lambda_{\min}$ be the condition number. Given $\epsilon,\eta\in(0,1)$, if the Lanczos iteration step $m$ and the query number $N$ satisfy
$m\geq\frac{1}{2\log(\rho_{1})}\cdot\log\left(K_{\rho_{1}}/\epsilon\right),$
and
$N\geq(24/\epsilon^{2})\cdot\left(\log(1+\kappa)\right)^{2}\cdot\log(2/\eta),$
where
$\rho_{1}=\frac{\sqrt{2\kappa+1}+1}{\sqrt{2\kappa+1}-1},K_{\rho_{1}}=\frac{8M_{\rho_{1}}}{\rho_{1}^{2}-\rho_{1}},M_{\rho_{1}}=5\log(2(\kappa+1)),$
then
$\mathbb{P}\Big{\{}\left|\log\det A-\left[\log\det A\right]_{L}\right|\leq\epsilon n\Big{\}}\geq 1-\eta.$
Proof 3.17.
Note that the parameter bound for the number of Rademacher distributed vectors $N$ is the same in the proof [33, Section 4.3]. Thus, the focus is on the lower bound of parameters related to the error between $(m+1)$-points Gauss quadrature $I_{m}$ and Riemann Stieljes integral $I$. In [33, Section 4.3], the authors commenced with the function $\log\left(\frac{t}{\lambda_{\max}+\lambda_{\min}}\right)$ that is analytic on $[\lambda_{\min},\lambda_{\max}]$. Here, one may simply focus on the analytic function $f(t)=\log(t)$ on $[\lambda_{\min}/(\lambda_{\max}+\lambda_{\min}),\lambda_{\max}/(\lambda_{\max}+\lambda_{\min})]$. One can further find a linear transformation
$h(t)=\left(\frac{\kappa-1}{2(\kappa+1)}\right)t+\frac{1}{2}$
to define
$g(t)=f\left(h(t)\right),$
such that $g$ is analytic on $[-1,1]$. The semimajor axis was chosen as $\alpha_{1}=(\kappa+1)/\kappa$ so the elliptical radius is
$\rho_{1}=\alpha_{1}+\sqrt{\alpha_{1}^{2}-1}=\frac{\sqrt{2\kappa+1}+1}{\sqrt{2\kappa+1}-1},$
and $g(t)$ is analytically continuable on $E_{\rho_{1}}$ with foci ±$\pm 1$. Note that the maximum of $g$ on the Bernstein ellipse $E_{\rho_{1}}$ equals the one of $f$ on the scaled Bernstein ellipse $E_{\rho_{1}}^{*}$. Further, the maximum value of $f$, i.e., $M_{\rho_{1}}$, is calculated on $E_{\rho_{1}}^{*}$ with foci $1/(\kappa+1)$ and $\kappa/(\kappa+1)$
$\max_{t\in E_{\rho_{1}}^{*}}|f(t)|$ $\leq\max_{t\in\partial_{E_{\rho_{1}}^{*}}}\sqrt{(\log|t|)^{2}+\pi^{2}}$
$=\sqrt{(\log|1/2\kappa|)^{2}+\pi^{2}}$
$\leq 5\log(2(\kappa+1))=M_{\rho_{1}},$
with the maximum obtained at $h(-\alpha_{1})=1/2\kappa$ according to the maximum modulus theorem [4, p. 175-178]. Based on Theorem 3.12, error between $I_{m}$ and $I$ has an upper bound
$|I-I_{m}|\leq\frac{4M_{\rho_{1}}}{(\rho_{1}^{2}-\rho_{1})}\rho_{1}^{-2m}.$
Then the absolute error bound between SLQ’s approximation and $N$-query Girard-Hutchinson estimator is
$|\operatorname{tr}_{N}(f(A))-\left[\log\det A\right]_{L}|=\frac{n}{N}\sum_{i=1}^{N}|I^{(i)}-I_{m}^{(i)}|=\frac{4nM_{\rho_{1}}}{(\rho_{1}^{2}-\rho_{1})\rho_{1}^{2m}}\leq\frac{\epsilon n}{2}.$
Therefore, one should take
$m\geq\frac{1}{2\log(\rho_{1})}\log\left(K_{\rho_{1}}/\epsilon\right),$
$K_{\rho_{1}}=\frac{8M_{\rho_{1}}}{\rho_{1}^{2}-\rho_{1}},M_{\rho_{1}}=5\log(2(\kappa+1)),$
to guarantee the probabilistic error bound stated in Theorem 3.16. Other parameters’ bounds are kept unchanged compared with Corollary 3.3.
Next, a probabilistic relative error bound for log determinant estimation problem would be derived. We follow the framework of [33, Theorem 4.1] and propose Theorem 3.18 under the assumption $\lambda_{\max}<1$ to give straight theoretical guidance on choosing parameters for SLQ that control the relative error under certain probability. This assumption is mild since for any SPD matrix with $\lambda_{\max}>1$, it can be simply scaled to have eigenvalues $\lambda\in(0,1)$.
Theorem 3.18.
Let $A\in\mathbb{R}^{n\times n}$ be a real symmetric positive definite matrix with eigenvalues in $[\lambda_{\min},\lambda_{\max}]$, $\lambda_{\max}<1$, and $\kappa=\lambda_{\max}/\lambda_{\min}$ be the condition number. Given $\epsilon,\eta\in(0,1)$, if the Lanczos iteration step $m$ and the query number $N$ satisfy
$m\geq\frac{1}{2\log(\rho_{2})}\cdot\log\left(\frac{K_{\rho_{2}}}{\epsilon\log\left(\sqrt[n]{\kappa}/\lambda_{\max}\right)}\right),$
and
$N\geq(24/\epsilon^{2})\cdot\log(2/\eta),$
where
$\rho_{2}=\frac{\lambda_{\max}+\sqrt{2\lambda_{\min}\lambda_{\max}-\lambda_{\min}^{2}}}{\lambda_{\max}-\lambda_{\min}},$
and
$K_{\rho_{2}}=\frac{8M_{\rho_{2}}}{\rho_{2}^{2}-\rho_{2}},M_{\rho_{2}}=\sqrt{\log\left(\frac{\lambda_{\min}}{2}\right)^{2}+\pi^{2}},$
then
$\mathbb{P}\Big{\{}\left|\log\det A-\left[\log\det A\right]_{L}\right|\leq\epsilon|\log\det A|\Big{\}}\geq 1-\eta.$
Proof 3.19.
Let
$h(t)=\left(\frac{\lambda_{\max}-\lambda_{\min}}{2}\right)t+\frac{\lambda_{\max}+\lambda_{\min}}{2}$
be a linear transformation and $g(t)=f(h(t))$, then $g(t)$ is analytic on $\left[-1,1\right]$. For function $f(t)=\log(t)$ that is analytic on $[\lambda_{\min},\lambda_{\max}]$, since function $g(t)$ has a singular value at $t=\mathring{t}=-(\lambda_{\max}+\lambda_{\min})/(\lambda_{\max}-\lambda_{\min})$, let $\alpha_{2}=\lambda_{\max}/(\lambda_{\max}-\lambda_{\min})<|\mathring{t}|$ as the semimajor axis length of $E_{\rho_{2}}$, with the elliptical radius
$\rho_{2}=\frac{\lambda_{\max}+\sqrt{2\lambda_{\min}\lambda_{\max}-\lambda_{\min}^{2}}}{\lambda_{\max}-\lambda_{\min}},$
then $g(t)$ is again analytically continuable in the open Bernstein ellipse $E_{\rho_{2}}$ with foci ±$\pm 1$.
In light of Theorem 3.12, one can derive
$\left|\operatorname{tr}_{N}(f(A))-\left[\log\det A\right]_{L}\right|\leq\frac{4nM_{\rho_{2}}}{(\rho_{2}^{2}-\rho_{2})\rho_{2}^{2m}},$ (17)
where $g(t)\leq M_{\rho_{2}}$ inside $E_{\rho_{2}}$.
Due to the maximum-modulus theorem [4, p. 175-178], $M_{\rho_{2}}$ is computed as
$\max_{t\in E_{\rho_{2}}}|g(t)|$ $=\max_{t\in\partial_{E_{\rho_{2}}}}\left|\log\left[\left(\frac{\lambda_{\max}-\lambda_{\min}}{2}\right)t+\left(\frac{\lambda_{\max}+\lambda_{\min}}{2}\right)\right]\right|$
$\leq\max_{t\in\partial_{E_{\rho_{2}}}}\sqrt{\left(\log\left|\left(\frac{\lambda_{\max}-\lambda_{\min}}{2}\right)t+\left(\frac{\lambda_{\max}+\lambda_{\min}}{2}\right)\right|\right)^{2}+\pi^{2}}$
$=\sqrt{\left(\log\left(\frac{\lambda_{\min}}{2}\right)\right)^{2}+\pi^{2}},$
because the maximum is obtained at $t=-\alpha=-\lambda_{\max}/(\lambda_{\max}-\lambda_{\min})$. In order to guarantee
$|\operatorname{tr}_{N}(f(A))-\left[\log\det A\right]_{L}|\leq\frac{\epsilon}{2}|\operatorname{tr}(f(A))|,$ (18)
let the $m$-related term in equation (17) be smaller than $\epsilon\log\left(\kappa/\lambda_{\max}^{n}\right)/2$, where
$\log\left(\kappa/\lambda_{\max}^{n}\right)=|(n-1)\log(\lambda_{\max})+\log(\lambda_{\min})|\leq|\operatorname{tr}(f(A))|.$
In this setting, let $K_{\rho_{2}}=8M_{\rho_{2}}/(\rho_{2}^{2}-\rho_{2})$ and the Lanczos step $m$ satisfies
$m\geq\frac{1}{2\log(\rho_{2})}\cdot\log\left(\frac{K_{\rho_{2}}}{\epsilon\log\left(\sqrt[n]{\kappa}/\lambda_{\max}\right)}\right),$
inequality (18) holds.
With the convergence analysis for the stochastic trace estimator first proposed by [1, Section 7] and further developed in [30, Theorem 1], when $N$, the number of queries, satisfies $N\geq(24/\epsilon^{2})\cdot\log(2/\eta)$, one may obtain
$\mathbb{P}\Big{\{}|\operatorname{tr}(f(A)-\operatorname{tr}_{N}(f(A))|\leq\frac{\epsilon}{2}|\operatorname{tr}(f(A))|\Big{\}}\geq 1-\eta.$
Thus,
$1-\eta\leq$$\mathbb{P}\left\{\left|\operatorname{tr}(f(A))-\operatorname{tr}_{N}(f(A))\right|\leq\frac{\epsilon}{2}|\operatorname{tr}(f(A))|\right\}$
$\leq$ $\mathbb{P}\Big{\{}\left|\operatorname{tr}(f(A))-\operatorname{tr}_{N}(f(A))\right|+\left|\operatorname{tr}_{N}(f(A))-\left[\log\det A\right]_{L}\right|$
$\quad\leq\frac{\epsilon}{2}|\operatorname{tr}(f(A))|+\frac{\epsilon}{2}\left|\operatorname{tr}(f(A))\right|\Big{\}}$
$\leq$ $\mathbb{P}\Big{\{}|\operatorname{tr}(f(A))-\left[\log\det A\right]_{L}|\leq\epsilon|\operatorname{tr}(f(A))|\Big{\}},$
that is
$\mathbb{P}\Big{\{}\left|\log\det A-\left[\log\det A\right]_{L}\right|\leq\epsilon\left|\log\det A\right|\Big{\}}\geq 1-\eta.$ (19)
Remark 3.20.
It should be pointed out that a larger semimajor axis implies a higher convergence rate of the Lanczos quadrature [33, Section 4.1]. In Theorem 3.18, semimajor axis $\alpha_{2}=\lambda_{\max}/(\lambda_{\max}-\lambda_{\min})$ is set greater than $\alpha_{1}=(\lambda_{\max}+\lambda_{\min})/\lambda_{\min}$ in Corollary 3.3 [33, Corollary 4.5]. On the other hand, $\alpha_{2}$ is smaller than $\alpha=(\lambda_{\max}+\lambda_{\min})/(\lambda_{\max}-\lambda_{\min})$ in [33, Theorem 4.1] but the log function would reach infinity near $t=-\alpha$, so $\alpha$ is invalid in this case.
4 Further discussion
In Theorem 3.16 and Theorem 3.18, both Lanczos quadrature and Girard-Hutchinson’s trace estimator set $\epsilon/2$ as the error bound. However, one may reallocate the error tolerance so that the total number of matrix vector multiplications (MVMs) is lessened. In brief, one should try to find a pair $(\alpha,\beta)$ such that
  1. 1.
    $|\operatorname{tr}_{N}(f(A))-\left[\log\det A\right]_{L}|\leq\frac{\epsilon}{\alpha}|\operatorname{tr}(f(A))|;$
  2. 2.
    $\mathbb{P}\left\{\left|\operatorname{tr}(f(A))-\operatorname{tr}_{N}(f(A))\right|\leq\frac{\epsilon}{\beta}|\operatorname{tr}(f(A))|\right\}\geq 1-\eta;$
  3. 3.
    $\frac{1}{\alpha}+\frac{1}{\beta}=1,\alpha>1,\beta>1,$
and equation (19) still holds. In this case, the steps of Lanczos iteration $m$ and the number of Rademacher vectors $N$ read
$m\geq\frac{1}{2\log(\rho_{2})}\log\left(\frac{K_{\rho_{2}}}{\epsilon\log\left(\sqrt[n]{\kappa}/\lambda_{\max}\right)}\right),$ (20)
and
$N\geq(6/\epsilon^{2})\cdot(\alpha/(\alpha-1))^{2}\cdot\log(2/\eta),$ (21)
where $K_{\rho_{2}}=\frac{4\alpha M_{\rho_{2}}}{\rho_{2}^{2}-\rho_{2}}$, $\rho_{2}$ and $M_{\rho_{2}}$ are defined in Theorem 3.18. Then from the perspective of minimizing MVMs with respect to $\alpha$, the problem can be viewed as
$\min_{\alpha>1}\,\log\left(\frac{4\alpha M_{\rho_{2}}}{\epsilon(\rho_{2}^{2}-\rho_{2})\log\left(\sqrt[n]{\kappa}/\lambda_{\max}\right)}\right)\left(\frac{\alpha}{\alpha-1}\right)^{2}.$
By introducing a constant $C=4M_{\rho_{2}}/(\epsilon(\rho_{2}^{2}-\rho_{2})\log\left(\sqrt[n]{\kappa}/\lambda_{\max}\right))$, the problem is simplified to
$\min_{\alpha>1}\,\log\left(C\alpha\right)\left(\alpha/(\alpha-1)\right)^{2}.$ (22)
The minimizer $\alpha^{*}$ of (22) then satisfies
$\alpha^{*}=2\log\alpha^{*}+2\log C+1,$ (23)
which is obtained through taking the derivative of the objective function and making it equal to 0. Numerical methods can be applied to solve equation (23). Theorem 4.1 summarizes the bounds of $m$ and $N$ which product is minimized through a reallocation of error.
Theorem 4.1.
Let $A\in\mathbb{R}^{n\times n}$ be a real symmetric positive definite matrix with eigenvalues in $[\lambda_{\min},\lambda_{\max}]$, $\lambda_{\max}<1$, and $\kappa=\lambda_{\max}/\lambda_{\min}$ be the condition number. Given $\epsilon,\eta\in(0,1)$, let $C=4M_{\rho_{2}}/(\epsilon(\rho_{2}^{2}-\rho_{2})\log\left(\sqrt[n]{\kappa}/\lambda_{\max}\right))$. If $\alpha^{*}=2\log\alpha^{*}+2\log C+1$, the Lanczos iteration step $m$ and the query number $N$ satisfy
$m\geq\frac{1}{2\log(\rho_{2})}\log\left(\frac{nK_{\rho_{2}}}{\epsilon\log\left(\sqrt[n]{\kappa}/\lambda_{\max}\right)}\right),$
and
$N\geq(6/\epsilon^{2})\cdot(\alpha^{*}/(\alpha^{*}-1))^{2}\cdot\log(2/\eta),$
where
$\rho_{2}=\frac{\lambda_{\max}+\sqrt{2\lambda_{\min}\lambda_{\max}-\lambda_{\min}^{2}}}{\lambda_{\max}-\lambda_{\min}},$
and
$K_{\rho_{2}}=\frac{4\alpha M_{\rho_{2}}}{\rho_{2}^{2}-\rho_{2}},M_{\rho_{2}}=\sqrt{\log\left(\frac{\lambda_{\min}}{2}\right)^{2}+\pi^{2}},$
then
$\mathbb{P}\Big{\{}\left|\log\det A-\left[\log\det A\right]_{L}\right|\leq\epsilon|\log\det A|\Big{\}}\geq 1-\eta.$
5 Experiments
A controllable experiment is designed to compare the numbers of MVMs required to guarantee the probabilistic error bound (19) in theory based on the results of Theorem 3.16, Theorem 3.18 and Theorem 4.1. We do not compare with the theorems proposed by Cortinovis and Kressner since their analyses are based on prior knowledge about matrices (e.g., $l_{2}$ norm and stable rank of $(\log(A)-\mathrm{diag}(\log(A))$) [6]. The test is carried out on four matrices with different decaying rates, such matrices can be kernel matrices corresponding to varying smoothness [34]. We also conduct experiments on two matrices that have practical applications [7]:
  1. (1)
    SPD matrix $A\in\mathbb{R}^{5000\times 5000}$ with eigenvalues $\lambda_{i}=0.99/i^{0.5},i=1,\ldots,5000$;
  2. (2)
    SPD matrix $A_{1}\in\mathbb{R}^{5000\times 5000}$ with eigenvalues $\lambda_{i}=0.99/i,i=1,\ldots,5000$;
  3. (3)
    SPD matrix $A_{2}\in\mathbb{R}^{5000\times 5000}$ with eigenvalues $\lambda_{i}=0.99/i^{2},i=1,\ldots,5000$;
  4. (4)
    SPD matrix $A_{3}\in\mathbb{R}^{5000\times 5000}$ with eigenvalues $\lambda_{i}=0.99/i^{3},i=1,\ldots,5000$;
  5. (5)
    SPD matrix from nd3k in [7];
  6. (6)
    SPD matrix from thermomec_TC in [7].
In the test, we set different levels of relative error $\epsilon^{*}=0.01,0.02,\ldots,0.2$, with fixed failure probability $\eta=0.1$. Note that Theorem 3.16 does not directly analyze the relative error of the log determinant approximation, but sets $\epsilon n$ as the absolute error bound. To ensure the comparability of results, $\epsilon$ in Theorem 3.16 should be replaced by
$\epsilon=\epsilon^{*}\log\det A/n.$ (24)
However, this move is somehow idealistic since the exact log determinant value in (24) is not known in advance. Hence, the numerical results of Theorem 3.16 in the comparison is biased to some extent. Despite the small defect, we may consider them as the benchmark.
For matrix $A$ with decaying rate $0.5$, Fig. 2a displays the numbers of MVMs required by different theorems to reach the varying probabilistic error bound. In terms of the other three synthetic matrices with different decaying rates, figures that depict the MVMs are much similar in shape so we would not plot them here. Fig. 2b shows that the error reallocation-based MVMs, as required by Theorem 4.1, would be smaller when eigenvalues decay slower. In Fig. 2c and Fig. 2d, one may also observe that the strategy of finding the suitable error allocation would reduce the approximations of log determinant for real matrices in the University of Florida sparse matrix collection [7].
Figure 2: Comparison of total MVMs suggested by Theorem 3.16, Theorem 3.18, Theorem 4.1 under failure probability $\eta=0.1$. (a) Results for $A\in\mathbb{R}^{5000\times 5000}$ with eigenvalues $\lambda_{i}=0.99/i^{0.5},i=1,\ldots,5000$; (b) Results of Theorem 4.1 for $A$, $A_{1}$, $A_{2}$ and $A_{3}\in\mathbb{R}^{5000\times 5000}$ that have different decaying rates $r=0.5,1,2,3$ with eigenvalues $\lambda_{i}=0.99/i^{r},i=1,\ldots,5000$; (c) Results for matrix from nd3k in [7]; (d) Results for matrix from thermomec_TC in [7].
6 Conclusion
In this paper, we reinvestigate the SLQ method for log determinant approximation problem. We articulate why the inconsistencies arise between [33] and [6]. It turns out that Ubaru, Chen and Saad’s analyses of SLQ is valid for symmetric quadrature nodes. But it seems that asymmetric cases arise more often in practice. An unnecessary scaling factor in the error of quadrature occurs in the proof of [33] when interval shifts between the physical domain and the reference domain. However such a tiny flaw doesn’t impact the decisive conclusion in [33]. It is based on such a framework that we develop analyses of SLQ for asymmetric quadrature nodes case. Such updated analyses will be required for our further study on the SLQ method [17]. Further work which details whether a Gauss quadrature rule is symmetric will be released in [16].
In this paper, we also apply an optimized error allocation technique that lessens the minimum requirements of total MVMs in theory. Experiment results demonstrate that such a technique is valid and might be of practical significance. Despite the advancements achieved in this field, a huge discrepancy persists between the theoretically determined number of MVMs and the actual number required, highlighting the need for further investigation to bridge this gap and optimize computational efficiency.
Appendix A Jacobi Iteration Matrices have Symmetric Eigenvalue Distribution
As first pointed out by [21, Section 3.7], matrices of form (10) have symmetric eigenvalue distribution, while Theorem 1.2.2 in [3] summarized the discussion.
Theorem A.1.
[3, Theorem 1.2.2] Let the singular value decomposition of $B\in\mathbb{C}^{m\times n}$ be $B=U\Sigma V^{H},$ where $\Sigma=\begin{bmatrix}\Sigma_{1}&O\\ O&O\end{bmatrix}$, $\Sigma_{1}=\mathrm{diag}(\sigma_{1},\sigma_{2},\cdots,\sigma_{r})$,
$U=\left[U_{1},U_{2}\right],U_{1}\in\mathbb{C}^{m\times r},U_{2}\in\mathbb{C}^{m\times(m-r)},V=\left[V_{1},V_{2}\right],V_{1}\in\mathbb{C}^{n\times r},V_{2}\in\mathbb{C}^{n\times(n-r)}.$
Then
$C=\begin{bmatrix}O&B\\ B^{H}&O\end{bmatrix}=P^{H}\begin{bmatrix}\Sigma_{1}&O&O&O\\ O&-\Sigma_{1}&O&O\\ O&O&O&O\\ O&O&O&O\end{bmatrix}P,$
where $P$ is unitary
$P=\frac{1}{\sqrt{2}}\begin{bmatrix}U_{1}&U_{1}&\sqrt{2}U_{2}&O\\ V_{1}&-V_{1}&O&\sqrt{2}V_{2}\end{bmatrix}^{H}.$
Hence the eigenvalues of $C$ are ±$\pm\sigma_{1},\pm\sigma_{2},\cdots,\pm\sigma_{r}$, and zero repeated $(m+n-2r)$ times, where $r=\mathrm{rank}(B)$.
Proof A.2.
One may directly check the matrix matrix product
$\frac{1}{2}\begin{bmatrix}U_{1}&U_{1}&\sqrt{2}U_{2}&O\\ V_{1}&-V_{1}&O&\sqrt{2}V_{2}\end{bmatrix}\begin{bmatrix}\Sigma_{1}&O&O&O\\ O&-\Sigma_{1}&O&O\\ O&O&O&O\\ O&O&O&O\end{bmatrix}\begin{bmatrix}U_{1}^{H}&V_{1}^{H}\\ U_{1}^{H}&-V_{1}^{H}\\ \sqrt{2}U_{2}^{H}&O\\ O&\sqrt{2}V_{2}^{H}\end{bmatrix}$
$=\frac{1}{2}\begin{bmatrix}U_{1}\Sigma_{1}&-U_{1}\Sigma_{1}&O\\ V_{1}\Sigma_{1}&V_{1}\Sigma_{1}&O\end{bmatrix}\begin{bmatrix}U_{1}^{H}&V_{1}^{H}\\ U_{1}^{H}&-V_{1}^{H}\\ \sqrt{2}U_{2}^{H}&O\\ O&\sqrt{2}V_{2}^{H}\end{bmatrix}$
$=\frac{1}{2}\begin{bmatrix}U_{1}\Sigma_{1}U_{1}^{H}-U_{1}\Sigma_{1}U_{1}^{H}&U_{1}\Sigma_{1}V_{1}^{H}+U_{1}\Sigma_{1}V_{1}^{H}\\ V_{1}\Sigma_{1}U_{1}^{H}+V_{1}\Sigma_{1}U_{1}^{H}&V_{1}\Sigma_{1}V_{1}^{H}-V_{1}\Sigma_{1}V_{1}^{H}\end{bmatrix}$
$=\begin{bmatrix}O&B\\ B^{H}&O\end{bmatrix}=C,$
which indicates that ±$\pm\sigma_{1},\pm\sigma_{2},\cdots,\pm\sigma_{r}$ are the eigenvalues of $C$, and zero repeated $(m+n-2r)$ times.
Further, it is trivial to deduce
Corollary A.3.
Let the singular value decomposition of $B\in\mathbb{C}^{m\times n}$ be $B=U\Sigma V^{H},$ where $\Sigma=\begin{bmatrix}\Sigma_{1}&O\\ O&O\end{bmatrix}$, $\Sigma_{1}=\mathrm{diag}(\sigma_{1},\sigma_{2},\cdots,\sigma_{r})$,
$U=\left[U_{1},U_{2}\right],U_{1}\in\mathbb{C}^{m\times r},U_{2}\in\mathbb{C}^{m\times(m-r)},V=\left[V_{1},V_{2}\right],V_{1}\in\mathbb{C}^{n\times r},V_{2}\in\mathbb{C}^{n\times(n-r)}.$
Then
$C=\begin{bmatrix}\gamma I_{m}&B\\ B^{H}&\gamma I_{n}\end{bmatrix}=P^{H}\begin{bmatrix}\gamma I_{r}+\Sigma_{1}&O&O&O\\ O&\gamma I_{r}-\Sigma_{1}&O&O\\ O&O&\gamma I_{m-r}&O\\ O&O&O&\gamma I_{n-r}\end{bmatrix}P,$
where $P$ is unitary
$P=\frac{1}{\sqrt{2}}\begin{bmatrix}U_{1}&U_{1}&\sqrt{2}U_{2}&O\\ V_{1}&-V_{1}&O&\sqrt{2}V_{2}\end{bmatrix}^{H}.$
Hence the eigenvalues of $C$ are $\gamma\pm\sigma_{1},\gamma\pm\sigma_{2},\cdots,\gamma\pm\sigma_{r}$, and $\gamma$ repeated $(m+n-2r)$ times, where $r=\mathrm{rank}(B)$.
Appendix B Tests on Theorem 3.6
We test the Ritz values’ locations in four cases. The first three cases study the reproducible matrix $A=H\Lambda H^{T}\in\mathbb{R}^{n\times n}$ with the diagonal matrix
$\Lambda=\mathrm{diag}(\lambda_{1},\cdots,\lambda_{n}),$
and the Householder matrix $H$,
$H=I_{n}-2(\mathbf{1}_{n}\mathbf{1}^{T}_{n}),$
where $I_{n}$ represents the identity matrix of size $n$ and $\mathbf{1}_{n}$ denotes the vector of $n$ ones. The eigenvalues stored in $\Lambda$ are predetermined and $n=50$ is fixed in our tests. The fourth case focuses on the nd3k matrix from [7].
  • Case 1: $\lambda_{i}=i/50$, $\boldsymbol{v}=\mathbf{1}_{50}/\sqrt{50}$;
  • Case 2: $\lambda_{i}=1/(51-i)$, $\boldsymbol{v}=\mathbf{1}_{50}/\sqrt{50}$;
  • Case 3: $\lambda_{i}=i/50$, $\boldsymbol{v}=(1,2,\cdots,50)^{T}/\|(1,2,\cdots,50)^{T}\|$.
  • Case 4: nd3k matrix, $\boldsymbol{v}=(1,\dots,1,-1,\dots,-1)^{T}/\sqrt{9000}\in\mathbb{R}^{9000}.$
Details of the 4 cases can be seen in Table 1.
$A$: symmetric eigenvalue distribution? $\boldsymbol{\mu}_{1}$: symmetric absolute values?
Case 1 Yes Yes
Case 2 No Yes
Case 3 Yes No
Case 4 No No
Table 1: Details of 4 cases
Note that the figure of $\mu(t)$ would be central symmetric about $(\bar{\lambda},\mu(\bar{\lambda}))$ if $A$ has symmetric eigenvalue distribution and $\boldsymbol{\mu}_{1}$ has symmetric absolute values. Figure 3 depicts the measure function $\mu(t)$ and the $10$ Ritz values for the three cases, which illustrates that Theorem 3.6 is valid to some extent.
(a) Case 1
(b) Case 2
(c) Case 3
(d) Case 4
Figure 3: Locations of Ritz values for three cases.